# The Monge optimal transport barycenter problem

### v1

@@ Abstract
A novel methodology is developed for the solution of the data-driven Monge optimal transport barycenter problem, where the pushforward condition is formulated in terms of the statistical independence between two sets of random variables: the factors $z$ and a transformed outcome $y$. Relaxing independence to the uncorrelation between all functions of $z$ and $y$ within suitable finite-dimensional spaces leads to an adversarial formulation, for which the adversarial strategy can be found in closed form through the first principal components of a small-dimensional matrix. The resulting pure minimization problem can be solved very efficiently through gradient descent driven flows in phase space. The methodology extends beyond scenarios where only discrete factors affect the outcome, to multivariate sets of both discrete and continuous factors, for which the corresponding barycenter problems have infinitely many marginals. Corollaries include a new framework for the solution of the Monge optimal transport problem, a procedure for the data-based simulation and estimation of conditional probability densities, and a nonparametric methodology for Bayesian inference.
@@ 1 Introduction
A central problem in the analysis of data is to estimate how a set of variables $x\in\mathcal{X}$, the outcome, depends on a set of covariates $z\in\mathcal{Z}$, the factors, a dependence that can be fully characterized by the conditional distribution $\rho(x|z)$. One seeks to extract from $n$ observed data pairs $\{x_{i},z_{i}\}$, either an evaluation procedure for $\rho$ itself or a procedure to draw samples $\{x_{j}^{*}\}$ from $\rho(x|z^{*})$ for any target value $z^{*}$. This is particularly challenging when $z$ includes continuous components, since any particular value $z^{*}$ has small probability of having appeared among the $\{z_{i}\}$, even less of having shown up in enough observational pairs to warrant a statistical analysis based on those pairs alone.
A data-driven methodology for the simulation of conditional distributions based on the [Monge] optimal transport barycenter problem (OTBP) seeks a map
@@
that removes from $x$ the variability that $z$ can explain, i.e. such that the random variables $y$ and $z$ are independent. In order not to remove any additional variability from $x$, we select the map $T$ that minimizes the expected value of a total “transportation” cost
$C=\mathbb{E}_{\pi}[c(x,T(x,z))],$
where $\pi$ is the joint distribution of $x$ and $z$. The pairwise cost function $c(x,y)$ quantifies the deformation of the data incurred by moving $x$ to $y$. This results in an OTBP of the form
$\min_{y=T(x,z)}\mathbb{E}_{\pi}\left[c\left(x,y\right)\right]\quad\hbox{s.t.}\quad y\perp\!\!\!\perp z,$
[…]
$x^{*}_{i}=T^{-1}\left(y_{i},z_{*}\right),\quad y_{i}=T\left(x_{i},z_{i}\right).$
In words, we remove from $x_{i}$ the variability attributable to $z$ having adopted the value $z_{i}$, and then restore that variability but with $z=z_{*}$. The variable $y$ represents the variability in $x$ that $z$ does not explain.
Other uses of the OTBP include he following:
@@
    In order to eliminate the effect of confounding variables $z$ from the data $x$, we simply move the $\{x_{i}\}$ to their counterpart in the barycenter, $\{y_{i}=T(x_{i},z_{i})\}$. Examples include the removal of batch effects, the consolidation of different data bases, where $z$ represents the data source and, more generally, the removal of the confounding effects of any set of variables $z$ that are not considered in the study under way. The removal of the effects of known factors $z$ helps identifying further sources of variability by investigating $y$, a version of $x$ cleaned of the effects of $z$.
@@
    The barycenter problem permits not only simulating but also estimating conditional densities, and therefore yields a mode-free, non-parametric data-based procedure for Bayesian inference: given a prior distribution $\gamma_{pr}(z)$, a set of sample pairs $\{x_{i},z_{i}\}$ drawn from an unknown joint distribution $\pi(x,z)$ and the observed current value of $x$, estimate the posterior distribution $\gamma_{pos}(z|x)$.
@@
    The optimal transport problem, a particular case of the OTBP with only two marginals, yields a natural horizontal distance among distributions. It also serves as a powerful tool for density estimation and sampling.
This article develops a novel, efficient methodology to solve the optimal transport barycenter problem, providing the capability to both simulate and estimate $\rho(x|z)$. Along the way, it clarifies the relation between the Monge optimal transport barycenter problem as posed above and the Wasserstein barycenter problem [3]. It also provides as corollaries new methodologies for the solution of Monge’s regular optimal transport problem and for model-free Bayesian inference.
The methodology’s central component is an adversarial formulation of the pushforward condition, where the independence between the random variables $z$ and $y$ is posed in terms of test functions. When these test functions are restricted to finite-dimensional inner-product spaces, the optimal adversarial strategy can be expressed in terms of the first principal components of a matrix, reducing the problem to a pure minimization, which can be solved very efficiently through a flow-based gradient-descent procedure. The corresponding optimal map $y=T(x,z)$ can be inverted in closed-form, which facilitates conditional density estimation and simulation. The closed form inversion formula itself has intrinsic value, as it extracts from the data natural factors $\{f^{k}(z)\}$ that encode the dependence of $x$ on $z$.
@@ 1.1 Relation to prior work
The Kantorovich –or Wasserstein– optimal transport barycenter problem was introduced in [4], defining the barycenter $\mu_{*}$ of a set of distributions $\{\mu_{i}\}$ as the minimizer of a weighted sum of the squared Wasserstein distances between the $\{\mu_{i}\}$ and $\mu_{*}$. At first sight, this problem appears to differ substantially from the Monge OTBP that we address in this article, which given a joint distribution $\pi(x,z)$ between the outcome $x$ and factors $z$ that are not necessarily discrete, seeks a cost-minimizing map $y=T(x,z)$ such that the resulting random variable $y$ is independent of $z$, a problem introduced in [1, 5]. Yet, as discussed in section 2, the Monge OTBP is equivalent to the Wasserstein barycenter problem extended to general factors $z$, when the solution of the latter is supported on $z$-dependent maps. The extension of the OTBP to a continuous covariate $z$ was studied in [6] in the context of its connection to the multimarginal optimal transport in the limit of infinitely many marginals.
There is a rich literature regarding the numerical solution of the OTBP, typically in their Kantorovich formulation. Within the data-driven problem alone, there is more than one way to classify the most popular approaches. We can first distinguish between discrete methods, which assume that the marginals at hand consist of convex linear combinations of Dirac delta functions, and continuous methods, which work instead under the hypothesis that smooth probability density functions underly the data at hand. The first category includes methods that leverage the Sinkhorn algorithm [7, 8, 9] and linear programming-based methods [10, 11]. Among the first algorithms to treat the problem in a continuous setting are [12, 13], both based on the dual of Kantorovich formulation.
@@
Those algorithms solving the continuous problem can be further divided into families. One criterion, functional to this work, regards the nature of the parametrization of the maps pushing forward each $\rho(x|z)$ to the barycenter. Most algorithms parametrize this map via a deep neural network with problem-dependent architecture. An example of this approach in [14, 15] uses Convex Neural Networks to parametrize a potential related to the optimal map. A different approach is the flow-based methodology adopted in [5] and inspired by [16, 17, 18]. Flow-based numerical solvers do not require neural networks or an a priori parametrization of the map, as they rely on the composition of infinitesimal elementary maps. They lead naturally to the adoption of gradient descent methods, more straightforward than saddle point optimizers, whose convergence is harder to characterize [19, 20].
While the time complexity of the cited solvers varies, a state-of-the-art solver does not really exist. Depending on the problem at hand, one approach should be favored over another. For instance, when dealing with images defined in high-dimensional spaces, a Sinkhorn-based approach with fast convergence may be more suitable than other approaches if one is not interested in resolving sharp details that are smoothed out by the entropic regularizer [21]. The complexity of the methodology that we proposed, which scales linearly with the number of sample points, is to our knowledge better than in all previous approaches.
Another connected work concerns not the OTBP but the related vector quantile regression [22]. This also transforms a random variable through a factor dependent map to another that is independent of the factor. Unlike the barycenter, however, the target distribution –the quantile– is fixed by the user, and the transformation proposed is linear and computed through linear programming. We believe that the connection between the two problems, the OTBP and quantile regression, is worth exploring, as it could provide new interpretations for both, as well as new applications, particularly in economics.
@@
Our approach focuses on the statistical analysis of data with factors $z$ that typically include continuous components, therefore requiring the solution of OTBP problems with infinitely many marginals. To the best of our knowledge, the work presented here and in [23, 2, 5] are the only ones dealing numerically with this aspect of the problem. Another major point distinguishing the work presented here from the existing literature is the interpretation of the pushforward condition that drives the samples underlying the marginals towards the barycenter. As in [5], we characterize the pushforward condition in terms of the statistical independence between two random variables: the cofactors $z$ and $y=T(x,z)$. This statistical characterization of the barycenter is critical for a number of applications, such as removal of variability and factor discovery [1, 2], and treatment effect estimation [24, 25] (See [26] for a similar characterization of independence through reproducing kernel Hilbert spaces.) Directly related to the statistical interpretation of the push-forward condition is the ability to solve numerically the barycenter problem for continuous factors and under costs different from the canonical Euclidean distance. While there is some literature dealing with more general costs (see for instance [27, 28]), to the best of our knowledge, the only alternative work on the solution of this problem in the continuous setting –i.e. with infinitely many marginals– is our own previous work in [23, 2, 5]. The formulations in those articles differ substantially from the current one: the first solved a minimax problem for a potential $\psi(y,z)$, extending the attributable component methodology of [29] beyond nonlinear regression, the second developed BaryNet, a network-based algorithm, and the third formulated the push-forward condition in terms of a test function of the form $F(y,z)=\rho(y|z)$, whose estimation through kernels has a computational cost that grows quadratically with the number of samples. By contrast, the current proposal, which is based on flows in phase space and formulates the pushforward condition in terms of the first principal components of a matrix, has a complexity that scales linearly with the number of observations. An additional property of the new methodology is that it automatically identifies nonlinear features $\{f^{l}(z)\}$ that encode the dependence of $x$ on $z$, effectively performing factor extraction.
@@ 1.2 Plan of the article
This article is structured as follows. Section 2 discusses the formulation of the Monge OTBP, relates it to an extension of the Wasserstein barycenter problem, and justifies its use to identify hidden sources of variability. Section 3 introduces an adversarial formulation of independence, relaxed to finite dimensional functional spaces provided with a suitable inner product. This gives rise to a rather compact formulation of the problem in terms of the singular values of a matrix, discussed in Section 4, and an efficient procedure for its minimization through gradient descent. Section 5 derives a closed-form expression for the inverse $x=X(y,z)$ of the map $y=T(x,z)$, mediated by extracted factors. Section 6 discusses various implementation aspects: the choice of functional spaces, the determination of the penalization parameter and of the learning rate, the termination criteria and the virtues of solving various barycenter problems in a row, each contributing to further explain the variability of the outcome $x$. It also discusses the algorithm’s complexity. Section 7 solves the regular optimal transport problem (in its Monge formulation) through a suitable reduction of the more general OTBP. This is used in Section 8 to perform regular and conditional density estimation. Section 9 illustrates the methodology through numerical examples. We first use synthetic data sets to demonstrate various aspects of the OTBP, such as the simulation and estimation of conditional densities, model-free Bayesian inference and the uncovering of hidden explanatory factors. Then we apply the procedure to real data sets related to weather and climate. Finally, Section 10 summarizes the procedure and suggests avenues for further improvement.
@@ 2 A Monge formulation of the optimal transport barycenter problem
Given a joint distribution
$\pi(x,z)=\rho(x|z)\ \gamma(z)$
between two sets of variables: the outcome $x\in\mathcal{X}$ and the covariates $z\in\mathcal{Z}$, we seek a map
$y=T(x,z)\in\mathcal{X}$
that removes from $x$ the variability that $z$ can explain, i.e. such that the random variables $y$ and $z$ are independent. We require that the space $\mathcal{X}$ have the structure of a smooth manifold, while the space $\mathcal{Z}$ can include both continuous and discrete components. We will further assume that $\rho(x|z)$ is absolutely continuous uniformly over $z$, vanishing on small subsets of $\mathcal{X}$. In order to remove from $x$ only the variability that $z$ can account for, we select the map $T$ that minimizes the expected value of the total transportation cost $C=\mathbb{E}_{\pi}[c(x,y)]$, where $c(x,y)$, an externally provided pairwise cost function, measures the deformation of the data incurred by moving $x$ to $y$. The canonical choice for $c$ is the squared distance
@@
The resulting problem reads
$\min_{y=T(x,z)}\mathbb{E}_{\pi}[c(x,y)]\quad\hbox{s.t.}\quad y\perp\!\!\!\perp z,$ (2)
[…]
@@
where $\Pi(\rho,\mu)$ is the set of joint distributions having $\rho$ and $\mu$ as marginals. The differences between (3) and (2) stem from their conceptual origin: while (3) extends the geometrical notion of barycenter to sets of distributions equipped with the Wasserstein distance, (2) uses a map $T$ to remove from $x$ any variability that $z$ can explain. It appears almost coincidental that there should be any connection between the two!
Yet the two problems relate as follows. The $i\in[1,\ldots,p]$ in (3) correspond to the $z\in\mathcal{Z}$ in (2), their weights $\{\lambda_{i}\}$ to the distribution $\gamma(z)$, and the $\{\mu_{i}\}$ to the $\rho(x|z)$. Thus the barycenter problem in (3) is restricted to discrete covariates $z$, which play the role of indexes for the distributions $\{\mu_{i}\}$. By contrast, the $z$ in (2) is a random variable of general type linked to $x$ through their joint distribution $\pi$. The well posedness of the barycenter problem with infinitely many marginals has been studied in [6] in connection to the multi-marginal optimal transport problem.
Extending Kantorovich’s relaxation of the optimal transport problem [3], (3) considers general couplings $\xi$ between the $\mu_{i}$ and $\mu$, while (2) extends Monge’s original formulation of optimal transport [30] to the barycenter problem, restricting attention to maps $T$ that push forward the $\rho(x|z)$ to $\mu$. These maps are central to the applications that motivate (2), as they are used both for conditional density simulation and estimation. Importantly, they turn $y=T(x,z)$ into a random variable that derives from $x$ and $z$, which leads to another critical distinction: while the argument of the minimization in (3) is the barycenter $\mu_{*}$ of the $\mu_{i}$, the formulation in (2) does not involve the barycenter at all! The fact that $y$ in this formulation is a random variable gives meaning to the alternative requirement of independence between $y$ and $z$.
That the two problems are much closer than they appear at first sight follows from the fact that classical work [31] has shown that, for smooth distributions and under quite general assumptions, the solution to Kantorovich’s formulation of the optimal transport problem also solves Monge’s, a results that has been extended to the OTBP in [4, 6]. We put together the connection between the two problems in the framework of this article through the three theorems that follow. First, a partial equivalence between the two formulations is given by the following theorem:
Theorem 1.
@@
If the minimizing couplings $\xi_{z}^{*}$ for the first problem are supported on maps $Q_{z}$,
$\xi_{z}^{*}(x,y)=\pi(x,\mathcal{Z})\ \delta\left(y-Q_{z}(x)\right),$
then
$T^{*}(x,z)=Q_{z}(x),\quad\hbox{and consequently }\ C_{M}=C_{K}\ \hbox{ and }\forall z\ \mu^{*}=T^{*}(:,z)\#\rho(:|z).$
@@
Since $\xi_{z}^{*}$ solves problem 1, $Q_{z}\#\rho(:|z)=\mu^{*}$, so the joint distribution $\Theta(y,z)$ satisfies
$\Theta(y,z)=\mu^{*}(y)\gamma(z),$
which implies that $y\perp\!\!\!\perp z$. If there existed another $T\neq Q_{z}$ pushing forward all $\rho(x|z)$ to a single distribution $\mu(y)$ at a cost $C_{M}<C_{K}$, then $\xi_{z}(x,y)=\pi(x,\mathcal{Z})\ \delta\left(y-T(x,z)\right)$ would solve problem 1 with a smaller cost than the optimal $\xi_{z}^{*}$, a contradiction.
If $\rho(x|z)$ vanishes on small sets uniformly over $z$, the optimal $\xi_{z}^{*}$ are indeed supported on maps, which also proves the existence of solutions to the Monge OTBP:
Theorem 2.
If the joint distribution $\pi(x,z)$ is absolutely continuous in $x$ uniformly over $z$, then the barycenter $\mu(y)$ of the corresponding $\rho(x|z)$ under the canonical cost in (1)) is also absolutely continuous and the optimal couplings $\xi_{z}^{*}$ are supported on maps.
This theorem extends to general covariates Theorem 5.1 in [4], where it is proved for discrete $z$’s. A similar result for continuous $z$ was proved in [6], see Corollary 3.3.3 and Theorem 4.2.5. An alternative proof, not included here for conciseness, builds on the fact that any small set $A\in\mathcal{X}$ on which a distribution $\mu$ does not vanish can be mollified so that the optimal transport between any smooth distribution $\rho$ and the mollified $\mu$ has a smaller total transportation cost (The bandwidth of the mollification relates to the degree of smoothness of $\rho$, hence the requirement of uniformity of the latter over $z$.)
We might conclude from theorems 1 and 2 that the Monge and [extended] Wasserstein barycenter problems are equivalent when applied to smooth distributions: after all, their unique solutions map to each other. Yet this equivalence applies to the problems’ solutions, not to their formulations. All the applications described in this article, as well as the very methodology proposed for solving the problem numerically, are strongly based on the map $y=T(x,z)$ and its inverse $X(y,z)$, both parameterized by $z$ and defining one of the random variables $(x,y)$ in terms of the other.
One such application of the Monge OTBP is to uncover hidden sources of variability by removing the effects of known factors. The relation between a random variable $x\in\mathcal{X}$ and known covariates $z\in\mathcal{Z}$ can be specified alternatively through the conditional distribution $\rho(x|z)$ and through a functional relation
$x=\phi(w,z),\quad w\in\mathcal{W},\quad w\sim\nu(:|z),$
where the random variable $w$ represents all additional causes of variability in $x$, which we either do not currently consider, cannot measure or are simply not aware of. The function $\phi$ and the distribution $\nu$ underlying $w$ determine $\rho(x|z)$ uniquely, but more than one pair $(\nu,\phi)$ can give rise to the same $\rho$. One special such pair is provided by the solution to the OTBP:
Theorem 3.
Given any joint distribution $\pi(x,z)$ that vanishes on small sets in $\mathcal{X}$ uniformly over $z$, the random variable $x$ can be written as
$x=X(y,z),$
where $y=T(x,z)\sim\mu$ is the solution to the barycenter problem for $\rho(x|z)$, so $X(:,z)=T^{-1}(:,z)$.

### v2

@@ Abstract
A novel methodology is developed for the solution of the data-driven Monge optimal transport barycenter problem, where the pushforward condition is formulated in terms of the statistical independence between two sets of random variables: the factors $z$ and a transformed outcome $y$. Relaxing independence to the uncorrelation between all functions of $z$ and $y$ within suitable inner-product spaces leads to an adversarial formulation, for which the adversarial strategy can be found in closed form through the first principal components of a matrix, and the resulting pure minimization problem can be solved efficiently through flows in phase space. The methodology extends beyond scenarios where only discrete factors affect the outcome, to multivariate sets of both discrete and continuous factors, for which the corresponding barycenter problems have infinitely many marginals. Corollaries include a new framework for the solution of the Monge optimal transport problem, a procedure for the data-based simulation and estimation of conditional probability densities, and a nonparametric methodology for Bayesian inference.
@@ 1 Introduction
A central problem in the analysis of data is to estimate how a set of variables $x\in\mathcal{X}$, the outcome, depends on a set of covariates $z\in\mathcal{Z}$, the factors, a dependence that can be fully characterized by the conditional distribution $\rho(x|z)$. One seeks to extract from $n$ observed data pairs $\{x_{i},z_{i}\}$, either an evaluation procedure for $\rho$ itself or a procedure to draw samples $\{x_{j}^{*}\}$ from $\rho(x|z_{*})$ for any target value $z_{*}$. This is particularly challenging when $z$ includes continuous components, since any particular value $z_{*}$ has small probability of having appeared among the $\{z_{i}\}$, even less of having shown up in enough observational pairs to warrant a statistical analysis based on those pairs alone.
This article develops a data-driven methodology for the estimation and simulation of conditional distributions based on the [Monge] optimal transport barycenter problem (OTBP), seeking a map
@@
that removes from $x$ the variability that $z$ can explain, i.e. such that the random variables $y$ and $z$ are independent. In order not to remove any additional variability from $x$, one selects the map $T$ that minimizes the expected value of a total transportation cost $C=\mathbb{E}_{\pi}[c(x,T(x,z))]$, where $\pi$ is the joint distribution of $x$ and $z$. The pairwise cost function $c(x,y)$ quantifies the deformation of the data incurred by moving $x$ to $y$. This results in an OTBP of the form
$\min_{y=T(x,z)}\mathbb{E}_{\pi}\left[c\left(x,y\right)\right]\quad\hbox{s.t.}\quad y\perp\!\!\!\perp z,$ (1)
where the symbol $\perp\!\!\!\perp$ stands for independence. We can use the solution to this problem to simulate the conditional distribution $\rho(x|z)$ for a target value $z=z_{*}$, extracting $n$ samples $\{x^{*}_{i}\}\sim\rho(x|z_{*})$ through
$x^{*}_{i}=T^{-1}\left(y_{i},z_{*}\right),\quad y_{i}=T\left(x_{i},z_{i}\right),$
where $T$ and its inverse $T^{-1}$ are regarded as maps between $x$ and $y$ parameterized by $z$. This procedure first removes from $x_{i}$ the variability attributable to $z=z_{i}$, and then restores that variability with $z=z_{*}$, while the variable $y$ stores the variability in $x$ that $z$ does not explain.
Other uses of the OTBP include the following:
@@
    In order to eliminate the effect of confounding variables $z$ from the data $x$, we simply move the $\{x_{i}\}$ to their counterpart in the barycenter, $\{y_{i}=T(x_{i},z_{i})\}$. Examples include the removal of batch effects, the consolidation of different data bases, where $z$ represents the data source and, more generally, the removal of the confounding effects of any set of variables $z$ that are not considered in the study under way.
@@
    The barycenter problem permits not only simulating but also estimating conditional densities (see section 8), and therefore yields a model-free, non-parametric data-based procedure for Bayesian inference: given a prior distribution $\gamma_{pr}(z)$, a set of sample pairs $\{x_{i},z_{i}\}$ drawn from an unknown joint distribution $\pi(x,z)$ and the observed current value of $x$, estimate the posterior distribution $\gamma_{pos}(z|x)$.
@@
    The optimal transport problem (OTP), a particular case of the OTBP with only two marginals, yields a natural horizontal distance among distributions. It also serves as a powerful tool for density estimation and sampling.
This article proposes an efficient methodology to solve the OTBP, providing the capability to both simulate and estimate $\rho(x|z)$. It also clarifies the relation between the statistical formulation of the Monge OTBP problem as posed above and the geometrically based Wasserstein barycenter problem [6]: while the latter addresses the barycenter’s distribution $\mu$, the former is centered on the underlying random variable $y$ and its relation to $x$ and $z$. Corollaries include new methodologies for the solution of Monge’s OTP and for model-free Bayesian inference.
The methodology’s central component is a new adversarial formulation of the pushforward condition, posed in terms of the independence between the random variables $z$ and $y$ and enforced through test functions. When these test functions are restricted to inner-product spaces, the optimal adversarial strategy can be expressed in terms of the first principal components of a matrix, reducing the problem to a pure minimization, which can be solved efficiently through a flow-based gradient-descent procedure. The corresponding optimal map $y=T(x,z)$ can be inverted in closed-form, which facilitates conditional density estimation and simulation. The closed form inversion formula extracts from the data natural factors $\{f^{k}(z)\}$ that encode the dependence of $x$ on $z$.
@@ 1.1 Relation to prior work
The Kantorovich –or Wasserstein– OTBP was introduced in [6], defining the barycenter $\mu_{*}$ of a set of distributions $\{\mu_{i}\}$ as the minimizer of a weighted sum of the squared Wasserstein distances between the $\{\mu_{i}\}$ and $\mu_{*}$. As discussed in section 2, the distribution underlying the solution $y$ to Monge’s OTBP agrees with the solution to the Wasserstein barycenter problem extended to general factors $z$, when the latter is supported on $z$-dependent maps. The extension of the OTBP to a continuous covariate $z$ was studied in [8] in the context of its connection to the multimarginal optimal transport in the limit of infinitely many marginals.
@@
There is a rich literature on the numerical solution of the OTBP, typically in their Kantorovich formulation. While discrete methods regard the marginals as convex linear combinations of Dirac delta functions, continuous methods assume that smooth probability density functions underly the data. The first category includes methods that leverage the Sinkhorn algorithm [9, 10, 11] and linear programming-based methods [12, 13]. Among the first algorithms to treat the problem in a continuous setting are [14, 15], both based on the dual of Kantorovich formulation. Most algorithms adopting the continuous apporach parameterize the maps pushing forward each $\rho(x|z)$ to the barycenter via a deep neural network with problem-dependent architecture. An example of this approach in [16] uses Convex Neural Networks to parametrize a potential related to the optimal map. A different approach is the flow-based methodology adopted in [7] and inspired by [17, 18, 19]. Flow-based numerical solvers do not require neural networks or any a priori parametrization of the map. They lead naturally to the adoption of gradient descent methods, more straightforward than saddle point optimizers, whose convergence is harder to characterize [20, 21]. Another connected work concerns not the OTBP but the related vector quantile regression [23]. This also transforms a random variable through a factor dependent map to another that is independent of the factor. Unlike the barycenter, however, the target distribution –the quantile– is fixed by the user, and the transformation proposed is linear and computed through linear programming.
@@
Our approach focuses on the statistical analysis of data with factors $z$ that typically include continuous components, therefore requiring the solution of OTBP problems with infinitely many marginals. A major point distinguishing the work presented here from the existing literature is the interpretation of the pushforward condition that drives the samples underlying the marginals towards the barycenter. As in [7], we characterize the pushforward condition in terms of the statistical independence between two random variables: the cofactors $z$ and $y=T(x,z)$. This statistical characterization of the barycenter is critical for a number of applications, such as removal of variability and factor discovery [3, 4], and treatment effect estimation [25, 26] (See [27] for a similar characterization of independence through reproducing kernel Hilbert spaces.) Directly related to the statistical interpretation of the push-forward condition is the ability to solve numerically the barycenter problem for continuous factors and under costs different from the canonical Euclidean distance. While there is some literature dealing with more general costs (see for instance [28, 29]), to the best of our knowledge, the only alternative work on the solution of this problem in the continuous setting –i.e. with infinitely many marginals– is our own previous work in [24, 4, 7]. The formulations in those articles differ substantially from the current one: the first solved a minimax problem for a potential $\psi(y,z)$, extending the attributable component methodology of [30] beyond nonlinear regression, the second developed BaryNet, a network-based algorithm, and the third formulated the push-forward condition in terms of a test function of the form $F(y,z)=\rho(y|z)$, whose estimation through kernels has a computational cost that grows quadratically with the number of samples. By contrast, the current proposal, which is based on flows in phase space and formulates the pushforward condition in terms of the first principal components of a matrix, has a complexity that scales linearly with the number of observations.
@@ 1.2 Plan of the article
This article is structured as follows. Section 2 discusses the formulation of the Monge OTBP, relates it to an extension of the Wasserstein barycenter problem, and justifies its use to identify hidden sources of variability. Section 3 introduces an adversarial formulation of independence, relaxed to finite dimensional inner-product functional spaces. This gives rise to a compact formulation of the problem in terms of the singular values of a matrix, discussed in Section 4, and an efficient procedure for its minimization through gradient descent. Section 5 derives a closed-form expression for the inverse $x=X(y,z)$ of the map $y=T(x,z)$, mediated by extracted factors. Section 6 discusses various implementation aspects: the choice of functional spaces, the determination of the penalization parameter and of the learning rate, the termination criteria and the solution of various barycenter problems in a row, each contributing to further explain the variability of the outcome $x$. It also discusses the algorithm’s complexity. Section 7 solves the regular [Monge] OTP through a suitable reduction of the more general OTBP. This is used in Section 8 to perform regular and conditional density estimation. Section 9 illustrates the methodology through numerical examples. We first use synthetic data sets to demonstrate various aspects of the OTBP, such as the simulation and estimation of conditional densities, model-free Bayesian inference and the uncovering of hidden explanatory factors. Then we apply the procedure to real data sets related to weather and climate. Finally, Section 10 summarizes the procedure and suggests avenues for further development.
@@ 2 A Monge formulation of the optimal transport barycenter problem
Given a joint distribution $\pi(x,z)=\rho(x|z)\ \gamma(z)$ between two sets of variables: the outcome $x\in\mathcal{X}$ and the covariates $z\in\mathcal{Z}$, we seek a map $y=T(x,z)\in\mathcal{X}$ that removes from $x$ the variability that $z$ can explain, i.e. such that the random variables $y$ and $z$ are independent. We require that the space $\mathcal{X}$ have the structure of a smooth manifold, while the space $\mathcal{Z}$ can include both continuous and discrete components. We will further assume that $\rho(x|z)$ is absolutely continuous for all $z$, vanishing on small subsets of $\mathcal{X}$. In order to remove from $x$ only the variability that $z$ can account for, we select the map $T$ that minimizes the expected value of the total transportation cost $C=\mathbb{E}_{\pi}[c(x,y)]$, where $c(x,y)$, an externally provided pairwise cost function, measures the deformation of the data incurred by moving $x$ to $y$. The canonical choice for $c$ is the squared distance
@@
At first sight, the resulting problem (1) looks quite different from the Wasserstein barycenter problem, introduced in [6], which reads
@@
where $\Pi(\rho,\mu)$ is the set of joint distributions having $\rho$ and $\mu$ as marginals. The differences between (3) and (1) stem from their conceptual origin: while (3) extends the geometrical notion of barycenter to sets of distributions equipped with the Wasserstein distance, (1) uses a map $T$ to remove from $x$ any variability that $z$ can explain. Yet the two problems relate as follows. The $i\in[1,\ldots,p]$ in (3) correspond to the $z\in\mathcal{Z}$ in (1), their weights $\{\lambda_{i}\}$ to the distribution $\gamma(z)$, and the $\{\mu_{i}\}$ to the $\rho(x|z)$. Thus the barycenter problem in (3) is restricted to discrete covariates $z$, which play the role of indexes for the distributions $\{\mu_{i}\}$. By contrast, the $z$ in (1) is a random variable of general type linked to $x$ through their joint distribution $\pi$. The well posedness of the barycenter problem with infinitely many marginals has been studied in [8] in connection to the multi-marginal OTP. Extending Kantorovich’s relaxation of the OTP [5], (3) considers general couplings $\xi$ between the $\mu_{i}$ and $\mu$, while (1) extends Monge’s original formulation [31] to the barycenter problem, restricting attention to maps $T$ that push forward the $\rho(x|z)$ to $\mu$. These maps are central to the applications that motivate (1), as they are used both for conditional density simulation and estimation. Importantly, they turn $y=T(x,z)$ into a random variable that derives from $x$ and $z$, which leads to another critical distinction: while the argument of the minimization in (3) is the barycenter $\mu_{*}$ of the $\mu_{i}$, the formulation in (1) does not involve the barycenter at all. The fact that $y$ in this formulation is a random variable gives meaning to the alternative requirement of independence between $y$ and $z$.
That the two problems are much closer than they appear at first sight follows from the fact that, as proved in [32], the solution to Kantorovich’s formulation of the OTP for smooth distributions and under quite general assumptions, also solves Monge’s, a results that has been extended to the OTBP in [6, 8]. We put together the connection between the two problems in the framework of this article through the discussion that follows. First, the following lemma shows that the two characterizations of the barycenter problem: one geometrical, in terms of the distribution with minimal Wasserstein distance from the marginals, and the other statistical, in terms of the independence between two random variables, are equivalent.
Lemma 1.
@@
If the minimizing couplings $\xi_{z}^{*}$ for the first problem are supported on maps, i.e., $\xi_{z}^{*}(x,y)=\pi(x,\mathcal{Z})\ \delta\left(y-Q_{z}(x)\right)$, then $T^{*}(x,z)=Q_{z}(x)$, and consequently $C_{M}=C_{K}$ and $\forall z\ \mu^{*}=T^{*}(:,z)\#\rho(:|z)$.
@@
Since $\xi_{z}^{*}$ solves problem 1, $Q_{z}\#\rho(:|z)=\mu^{*}$, so the joint distribution satisfies $\Theta(y,z)=\mu^{*}(y)\gamma(z)$, which implies that $y\perp\!\!\!\perp z$. If there existed another $T\neq Q_{z}$ pushing forward all $\rho(x|z)$ to a single distribution $\mu(y)$ at a cost $C_{M}<C_{K}$, then $\xi_{z}(x,y)=\pi(x,\mathcal{Z})\ \delta\left(y-T(x,z)\right)$ would solve problem 1 with a smaller cost than the optimal $\xi_{z}^{*}$, a contradiction.
If the joint distribution $\pi(x,z)$ is absolutely continuous in $x$ for a set of $z\in Z$ of nonzero measure, then the barycenter $\mu(y)$ of the corresponding $\rho(x|z)$ under the canonical cost in (2)) is also absolutely continuous and the optimal couplings $\xi_{z}^{*}$ are indeed supported on maps. This result was proved for discrete $z$’s in [6], Theorem 5.1, and extended to continuous $z$ in [8], see Corollary 3.3.3 and Theorem 4.2.5. We might conclude that the Monge and [extended] Wasserstein barycenter problems are equivalent when applied to smooth distributions: after all, their unique solutions map to each other. Yet this equivalence applies to the problems’ solutions, not to their formulations. All the applications described in this article, as well as the methodology proposed for solving the problem numerically, are strongly based on the map $y=T(x,z)$ and its inverse $X(y,z)$, both parameterized by $z$ and defining the random variables $(x,y)$ in terms of each other.
The relation between a random variable $x\in\mathcal{X}$ and known covariates $z\in\mathcal{Z}$ can be specified alternatively through the conditional distribution $\rho(x|z)$ and through a functional relation $x=\phi(w,z)$, $w\in\mathcal{W}$, $w\sim\nu(:|z)$, where the random variable $w$ represents all additional causes of variability in $x$, which we either do not currently consider, cannot measure or are simply not aware of. The function $\phi$ and the distribution $\nu$ underlying $w$ determine $\rho(x|z)$ uniquely, but more than one pair $(\nu,\phi)$ can give rise to the same $\rho$. One special pair is provided by the solution to the OTBP:
Lemma 2.
Given any joint distribution $\pi(x,z)$ that vanishes on small sets in $\mathcal{X}$ for all $z$, the random variable $x$ can be written as $x=X(y,z)$, where $y=T(x,z)\sim\mu$ is the solution to the barycenter problem for $\rho(x|z)$, so $X(:,z)=T^{-1}(:,z)$.